1. You normally drive on the freeway between Lopez, Quezon and San Pablo, Laguna at an


average speed of 105 km/hr, and the trip takes 2 hours and 20 minutes. On a Friday


afternoon, however, heavy traffic slows you down and you drive the same distance at an


average speed of only 70 km/hr. How much longer does the trip take?

1
Expert's answer
2022-04-17T17:05:59-0400

The distance can be found as follows:


"d = v_1 t_1"

where "v_1=105km\/h", "t_1 = 2h\space 20min=7/3h". The time in the second case is:


"t_2 = \dfrac{d}{v_2} = \dfrac{v_1t_1}{v_2}"

where "v_2 = 70 km\/h". Thus, obtain:


"t_2 = \dfrac{105km/h\cdot 7/3h}{70km/h}=3.5h=3h\space 30 min"

Answer. It will take 3h 30min.